# RapidPoseTriangulation: Multi-view Multi-person Whole-body Human Pose Triangulation in a Millisecond

### v2

@@ 1 Introduction
However, marker-less pose estimation presents its own set of challenges. Occlusions, varying lighting conditions, and the complexity of human movements can all impact the accuracy and reliability of these systems. To address external and self-occlusion issues, and to some extent also varying lighting, multi-view approaches have become more common. By capturing the scene from multiple angles simultaneously, these systems can overcome many of the limitations that would occur with single-view methods.
@@ 2 Related Work
On the side of the learning-based methods, VoxelPose [20] was one of the first concepts, building upon the research of Iskakov et al. [11] and extending it to multi-person estimations. This method projects joint heatmaps from 2D images into a 3D voxelized space, then first estimates a central point for each individual, which is subsequently utilized to generate and extract a focused cube surrounding each person. Afterward, a secondary neural network computes the joint positions. Faster-VoxelPose [23] enhanced this technique by restructuring the 3D voxel network into several 2D and 1D projections, thereby boosting efficiency. TesseTrack [15] and TEMPO [6] incorporated a temporal dimension into the voxel space to refine poses across sequential frames. Alternative approaches such as PRGnet [22] employ a graph-based methodology or directly infer 3D poses from 2D features, like in MvP [21]. SelfPose3d [18] adopts the framework of VoxelPose and trains both 2D and 3D networks in a self-supervised manner using randomly augmented 2D poses.
@@ 3.1 Algorithm
Figure 2: Obtaining 3D proposals. The process starts with 2D detections for each view (a, images from chi3d [9]). Then all possible pairs between the views are created. In this case this leads to six pairs from the two detections above with the three below. The filter step (2) with previous poses is skipped here. The core joints of all pairs are triangulated into 3D proposals (b, steps (3,4)). Then they are reprojected into the 2D views (c, step (6)), and a distance-based error to the original 2D poses (visualized in black in c) is calculated. As can be seen in the image, the green and pink proposals clearly do not match to their original 2D poses, and get a very high error. The yellow and light-blue poses resulted from the flipped (man with woman) pairs and also have a notable error. All pairs with errors above a threshold are dropped in step (8). Only the remaining red and dark-blue pairs with low enough errors are used for the further steps (9-12).
@@
In step (1) for each 2D pose, all possible pairs with poses from other views are created, since no association to real persons is known yet. In the next step (2), the pairs are filtered using the previous 3D poses, if available. For this, the 3D poses from the last frame are projected into the 2D views and a pixel-based distance threshold is used. The idea is to reduce the total number of pairs that need to be triangulated. In a pair of 2D persons in two different views, either both parts can match to one of the reprojected 3D poses, only one part is matched, or none of them. If both pairs match, it is likely a valid pair belonging to that 3D person, and it is kept. If only one part matches, a match is unlikely, because the 3D pose should be matchable in both views, so the second part is likely a pose from a different 3D person, and the pair is dropped. If none is matching, it is likely that a new 3D person is present in the current frame, and the pair is kept. In step (3) the core joints are extracted, which are the shoulders, hips, elbows, wrists, knees, and ankles. They are enough to associate the following 3D proposals to a person, and this reduces the computational complexity in the next steps. In step (4) each pair is triangulated to a 3D proposal (see Figure 2(b)). In the following step (5) clearly invalid proposals outside the observed room are dropped. In step (6) the remaining proposals are reprojected into the 2D views. Step (7) calculates a pixel-distance-based reprojection error to the pair’s original 2D poses, and also prunes clearly invalid limbs. In step (8) pairs/proposals with a large error are dropped, as they are considered invalid matches. In step (9) the remaining 3D proposals are grouped in 3D space. Each of the 2D view pairs created a 3D proposal, and if they are close in 3D space, they likely belong to the same person, and are therefore grouped together. Grouping in 3D makes use of physical constraints, which allows a simpler and faster implementation. In step (10) the remaining pairs, which normally are less than in step (4), are triangulated again, but this time with all joints. In step (11) for each group the calculated 3D proposals are merged into a single 3D pose. To filter outliers, first, for each joint an average location is calculated, and then for each proposal, the distance of this joint to the average is calculated. If the distance is too great, the joint is dropped. From the remaining joint proposals, the top-k closest to the average are selected, to drop outliers again, and the final joint location is calculated as the average of those. In general, this late fusion step allows for stronger outlier filtering, because multiple proposals can be used, but also works if a person is only visible in two views. In step (12) clearly invalid persons, with too few keypoints, that are too small or large, or outside the room, are dropped, and missing joints are filled with their next neighbors. The general idea behind the algorithmic structure is to use only the minimal necessary information at each step, drop bad proposals early on, and move expensive operations, like the full-body keypoint calculations, to later steps.
@@ 3.3 Performance
Pair filtering (2) 8
Triangulate and Score (3-7) 27
Grouping (9) 6
Triangulate and Score (10) 27
@@
Post-process (12) 8
Total 111
@@ 4.2 Additional datasets
The first one is the chi3d [9] dataset, which shows two persons interacting with each other. It has the challenge that the persons are often very close to each other, which makes it hard to distinguish between them. The setup is similar to the one of human36m and also contains four cameras. The results in Table 8 show that generally most persons are detected, with a similar performance as in human36m. Regarding RapidPoseTriangulation, it could be seen that in a few cases in which the 2D pose estimation struggled to correctly detect the joints between very close person, the 3D triangulation did not get enough inputs to correctly triangulate the joints and could miss a person. One could reduce the reprojection matching threshold to better handle such problematic 2D detections, but this would also increase the number of false positives. The authors of CloseMoCap [16] also reported very good transfer results on this dataset, reaching a PCK@50 of $94.3$ on average. That would be better than any PCK@100 score reached here. They also evaluated VoxelPose, Faster-VoxelPose and mvpose on the same dataset and reached much better results for them as well. Besides that, they reported per-joint errors on a subset of the dataset and found that the hip joints had the best results, whereas here the hip joints are among the worst results. In the labels, it though can be seen that the hips are positioned higher on the body than in the COCO skeleton definition. So in conclusion the mismatch is likely caused by parts of the evaluation somewhere in CloseMoCap, but even though it was promised in its paper, the source-code was not published, so this assumption can not be verified.

### v3

@@ 1 Introduction
However, marker-less pose estimation, especially with only a single viewpoint, presents its own set of challenges. Occlusions, varying lighting conditions, and the complexity of human movements can all impact the accuracy and reliability of these systems. To address external and self-occlusion issues, and to some extent also varying lighting, multi-view approaches have become more common. By capturing the scene from multiple angles simultaneously, these systems can overcome many of those limitations.
@@ 2 Related Work
On the side of the learning-based methods, VoxelPose [20] was one of the first concepts, building upon the research of Iskakov et al. [11] and extending it to multi-person estimations. Its method projects joint heatmaps from 2D images into a 3D voxelized space, estimates a central point for each individual using a convolution network, extracts a focused cube surrounding each person center, and computes the joint positions using a second net. Faster-VoxelPose [23] enhanced this technique by restructuring the 3D voxel network into several 2D and 1D projections, thereby boosting efficiency. TesseTrack [15] and TEMPO [6] incorporated a temporal dimension into the voxel space to refine poses across sequential frames. Alternative approaches such as PRGnet [22] employ a graph-based methodology or directly infer 3D poses from 2D features, like in MvP [21]. SelfPose3d [18] adopts the framework of VoxelPose and trains both 2D and 3D networks in a self-supervised manner using randomly augmented 2D poses.
@@ 3.1 Algorithm
  13. 13.
    Optional: Assign persons to tracks over time
  14. 14.
    Optional: Clip joint movements to a maximum speed
@@
Figure 2: Obtaining 3D proposals. The process starts with 2D detections for each view (a, images from chi3d [9], note the blue colored false positive). Then, in step (1), all possible pairs between the views are created. In this case this leads to six pairs from the two detections above with the three below. The core joints of all pairs are triangulated into 3D proposals (b, steps (3,4)). Then they are reprojected into the 2D views (c, step (6)), and a distance-based error to the original 2D poses (visualized in black in c) is calculated. As can be seen in the image, the green and pink proposals clearly do not match to their original 2D poses, and get a very high error. The yellow and light-blue poses resulted from the flipped (man with woman) pairs and also have a notable error. All pairs with errors above a threshold are dropped in step (8). Only the remaining red and dark-blue pairs with low enough errors are used for the further steps (9-12).
@@
In step (1) for each 2D pose, all possible pairs with poses from other views are created, since no association to real persons is known yet. In the next step (2), the pairs are filtered using the previous 3D poses, if available. For this, the 3D poses from the last frame are projected into the 2D views and a pixel-based distance threshold is used. The idea is to reduce the total number of pairs that need to be triangulated. In a pair of 2D persons in two different views, either both parts can match to one of the reprojected 3D poses, only one part is matched, or none of them. If both pairs match, it is likely a valid pair belonging to that 3D person, and it is kept. If only one part matches, a match is unlikely, because the 3D pose should be matchable in both views, so the second part is likely a pose from a different 3D person, and the pair is dropped. If none is matching, it is likely that a new 3D person is present in the current frame, and the pair is kept. In step (3) the core joints are extracted, which are the shoulders, hips, elbows, wrists, knees, and ankles. They are enough to associate the following 3D proposals to a person, and this reduces the computational complexity in the next steps. In step (4) each pair is triangulated to a 3D proposal (see Figure 2(b)). In the following step (5) clearly invalid proposals outside the observed room are dropped. In step (6) the remaining proposals are reprojected into the 2D views. Step (7) calculates a pixel-distance-based reprojection error to the pair’s original 2D poses, and also prunes clearly invalid limbs. In step (8) pairs/proposals with a large error are dropped, as they are considered invalid matches. In step (9) the remaining 3D proposals are grouped in 3D space. Each of the 2D view pairs created a 3D proposal, and if they are close in 3D space, they likely belong to the same person, and are therefore grouped together. Grouping in 3D makes use of physical constraints, which allows a simpler and faster implementation. In step (10) the remaining pairs, which normally are less than in step (4), are triangulated again, but this time with all joints. In step (11) for each group the calculated 3D proposals are merged into a single 3D pose. To filter outliers, first for each joint an average location is calculated, and then for each proposal, the distance of this joint to the average is calculated. If the distance is too great, the joint is dropped. From the remaining joint proposals, the top-k closest to the average are selected, to drop outliers again, and the final joint location is calculated as the average of those. In general, this late fusion step allows for stronger outlier filtering, because multiple proposals can be used, but also works if a person is only visible in two views. In step (12) clearly invalid persons, with too few keypoints, that are too small or large, or outside the room, are dropped, and missing joints are filled with their next neighbors. The general idea behind the algorithmic structure is to use only the minimal necessary information at each step, drop bad proposals early on, and move expensive operations, like the full-body keypoint calculations, to later steps.
Optionally, in steps (13-14) the results can be tracked and smoothed over time. A simple distance-based threshold is used to assign poses to existing tracks. Just clipping the joint movements to a maximum speed was preferred to more complex smoothing filters, because it does not introduce response latency to sudden direction changes, which could hurt real-time position accuracy.
@@ 3.3 Performance
Pair filtering (2) 10
Triangulate and Score (3-7) 28
Grouping (9) 7
Triangulate and Score (10) 28
@@
Post-process (12) 7
Tracking (13-14) 1
Total 115
@@ 4.2 Additional datasets
The first one is the chi3d [9] dataset, which shows two persons interacting with each other. It has the challenge that the persons are often very close to each other, which makes it hard to distinguish between them. The setup is similar to human36m and also contains four cameras. The results in Table 8 show that generally most persons are detected, with a similar performance as in human36m. Regarding RapidPoseTriangulation, it could be seen that in some cases the 2D pose estimation struggled to correctly detect the joints between very close persons, and the 3D triangulation then did not get enough inputs to correctly triangulate all joints and could miss a person. One could reduce the reprojection matching threshold to better handle such problematic 2D detections, but at the cost of increasing the number of false positives, or alternatively use the tracking information to fill-up missed persons with their last positions. The authors of CloseMoCap [16] also reported very good transfer results on this dataset, reaching a PCK@50 of $94.3$ on average. They also evaluated VoxelPose, Faster-VoxelPose and mvpose on the same dataset and reached much better results for them as well. Besides that, they reported per-joint errors on a subset of the dataset and found that the hip joints had the best results, whereas here the hip joints are among the worst results. In the labels, it though can be seen that the hips are positioned higher on the body than in the COCO skeleton definition. So in conclusion the mismatch is likely caused by parts of the evaluation, or the use of different labels, somewhere in CloseMoCap, but because their source-code was not published, this assumption can not be verified.